In 2007 Person A's stock portfolio decreased in value by 20%. In 2008 his stock portfolio decreased in value by another 40%. What percentage gain is required in 2009 for person A's stock portfolio to return to the value it had at the beginning of 2007? round the answer to nearest tenth as needed
Answer by stanbon(75887)

In 2007 Person A's stock portfolio decreased in value by 20%. In 2008 his stock portfolio decreased in value by another 40%. What percentage gain is required in 2009 for person A's stock portfolio to return to the value it had at the beginning of 2007? round the answer to nearest tenth as needed
--------
2007 amount: x
end of 2007 amount : 0.80x
end of 2008 amount : 0.6*0.8x = 0.48x
---
Solve for "k": k(0.48x) = x
---
k = x/(0.48x) = 1/0.48 = 2.083
----
His 2008 amount must increase by 208.3%
=======================================
Cheers,
Stan H.